OpenStudy (anonymous):

b/7 = 7 show all work One step equation

OpenStudy (anonymous):

multiply both sides by 7.

OpenStudy (anonymous):

can you do it forme?

OpenStudy (ash2326):

We have \[\frac{b}{7}=7\] Multiply both sides by 7 \[\frac{b}{7} \times 7=7 \times 7\] \[\frac{b}{\cancel 7} \times \cancel 7=7 \times 7\] We get \[b=49\]

OpenStudy (anonymous):